# Local Advice and Local Decompression

### v1

@@ 6. $\Delta$-coloring of $\Delta$-colorable graphs
The remaining part of this section is dedicated to proving Theorem 1. The formal definition of advice schema and composability are presented in Definition 2 and Definition 4, respectively. To recall, in order to prove that a composable schema exists, for any given $\gamma\geq\gamma_{0}$ (and hence in our case a large-enough constant), and for any given $\alpha\geq A(c,\gamma)$, we need to provide a schema of distributing advice bits to the vertices of the graph such that, in any ball of radius $\alpha$ there are at most $\gamma_{0}$ bit-holding nodes, and bit-holding nodes must have at most $c\alpha/\gamma^{3}$ bits each. Moreover, given such a bit assignment, there must exist an algorithm that can solve $\Delta$-coloring in $\max(O(\alpha^{2}\log\Delta),O(\alpha\log\Delta\cdot\sqrt{\Delta\log\Delta})$ time.
@@
  2. (2)
    Then, the algorithm reduces the number of colors to $\Delta+1$, by using a standard distributed coloring algorithm, which runs in $O(\sqrt{\Delta\log\Delta})$ rounds when an initial $O(\Delta^{2})$-coloring is given (Fraigniaud et al., 2016; Barenboim et al., 2022; Maus and Tonoyan, 2022).
  3. (3)
    Finally, the algorithm reduces the number of colors from $\Delta+1$ to $\Delta$, inspired by distributed algorithms from (Panconesi and Srinivasan, 1992; Ghaffari et al., 2021), leveraging the advice bits to terminate in just $O(\log\Delta\cdot\sqrt{\Delta\log\Delta})$ rounds.
The second step is a simple application of distributed algorithms with no advice, while the first and third algorithms rely on some given advice. Below, we formally state the problems to be solved in steps 1 and 3, and then we design composable schemas for them. Later, we will compose the schemas into a single one by applying Lemma 1, obtaining Theorem 1.
6.1. $O(\Delta^{2})$-vertex coloring
@@
Problem 1 ($O(\Delta^{2})$-coloring).
@@
The algorithm roughly follows the approach that can be used to color a graph via a network decomposition, which can be described as follows. Let us consider a partition of vertices into disjoint clusters, each cluster having only $\Delta^{O(\alpha^{2}\log\Delta)}$ adjacent clusters and a small diameter. In order to find a proper $\Delta^{O(\alpha^{2}\log\Delta)}$ coloring, we can use a proper $\Delta+1$ coloring of vertices in each cluster, and a proper $\Delta^{O(\alpha^{2}\log\Delta)}$ coloring of the cluster graph (i.e., a coloring of the graph where each cluster is contracted to a point, or, in other words, an assignment of colors to clusters such that adjacent clusters have different colors). The color of a vertex $v$ in a cluster $C$ is the tuple ((((color of $v$ in $C$, color of $C)$. The first part of the tuple ensures that neighboring vertices in the same cluster have different colors, and the second part ensures that vertices in two adjacent clusters have different colors. The total number of colors we use is still $\Delta^{O(\alpha^{2}\log\Delta)}$. The number of colors we use can be then reduced to $O(\Delta^{2})$ by using a few rounds of Linial’s coloring algorithm (Linial, 1992).
Lemma 0.
@@
(Linial, 1992) Let us consider a graph with maximum degree $\Delta$ and some coloring of its vertices with colors $1,2,\dots,c$. There exists $2$ round distributed LOCAL algorithm that finds a proper $O(\Delta^{2})$ coloring, if $c\in\operatorname{poly}(\Delta)$, or an $O(\Delta^{2}\log c)$ coloring otherwise.
@@
Given that the total number of colors before we start using the reduction described in Lemma 4 is $\Delta^{O(\alpha^{2}\log\Delta)}$, we need $O(\log^{*}\alpha)$ rounds to reduce the number of colors down to $O(\Delta^{2})$.
@@
In the remaining part of the proof of Lemma 3 we show that we can define an adequate clustering, a distribution of advice bits, and an algorithm that using these advice bits finds a proper $\Delta^{O(\alpha^{2}\log\Delta)}$ coloring of the cluster graph.
@@
In the following, the degree of a cluster is the number of edges with exactly one endpoint in the cluster. The idea of the advice schema of this section is similar to the idea behind the schema for graphs with sub-exponential growth. Essentially, we show that we can define a clustering with clusters that have sufficiently many internal vertices to encode the color of the cluster. However, to provide such a clustering, we rely on a slightly different idea, namely: we split all clusters into $O(\alpha^{2}\log\Delta)$ buckets, and let the $i$th bucket contain every cluster whose degree is in the interval $[\Delta^{i-1},\Delta^{i})$. Each bucket receives its own disjoint palette of colors; the $i$th bucket receives a palette of size $\Delta^{i}$. The clusters from the $i$th bucket will be colored only with those colors, and as such the number of bits describing such a color is at most $(i+1)\log\Delta$. For any $\Delta>1$, the total number of colors in bucket $i$ and all smaller buckets is $\sum_{j=1}^{i}\Delta^{j}<\Delta^{i+1}$.
@@
To define the clustering, we proceed as follows. Let $r=100\alpha^{2}\log\Delta$. We pick an $(r,r)$-ruling set $I$ from $G$. Then, we assign each vertex from $G$ to the closest vertex from $I$ (breaking ties in an arbitrary consistent manner, e.g., by joining the cluster of the smallest ID neighbor). That concludes the first part of our construction. Currently, we have clusters satisfying that all nodes that are within distance $r/2-1=50\alpha^{2}\log\Delta-1$ from a cluster center $v$ are guaranteed to belong to the cluster of $v$, and such that the degree of the clusters is bounded from above by $\Delta^{r+1}$. However, for our advice schema, we need the clusters with degrees larger than $\Delta^{10\alpha}$ to have at least $2100\alpha^{2}\log^{2}\Delta+$ 2 internal vertices that are at least $\alpha$ hops away from the border of the cluster.
@@
We deem a cluster to be broken when it has degree at least $\Delta^{10\alpha}$ but does not have at least $\Delta^{9\alpha-2}$ vertices of distance at least $\alpha$ from its border. In the refining process, we consider all broken clusters one by one, and we ‘repair’ each broken cluster by destroying its neighbors, and possibly destroying other nearby clusters. Whenever we consider a broken cluster $B$ that is still in $C$, we remove from $C$ all other clusters ($\neq B$) that contain at least one vertex whose distance from the border of $B$ is smaller than $12\alpha+$ 1. Thus, the set $I^{\prime}$ still contains the center $i_{B}$ of said broken cluster, but has lost one or more other cluster centers.
@@
Proof.
For any broken cluster $B$ that is not removed from $C$, the refining process removed all clusters (in particular all high-degree clusters) with vertices at a distance at most $22\alpha+$ 2 from the border of $B$. As such, the re-partition step assigns to $B$ all vertices that are at distance at most $\alpha$ from B’s border, as any other cluster is at least $\alpha+1$ hops away, and $B$ is at most $\alpha$ hops away. Thus, all vertices that were originally in $B$ are now at least $\alpha$ hops away from their new border. In conclusion, $B$ contains at least $\Delta^{10\alpha}$ nodes that are at least $\alpha$ hops away from their new border. ∎
@@
Clusters with degree at most $\Delta^{10\alpha}$
If the degree of a cluster is smaller than $\Delta^{10\alpha}$, we need $10\alpha\log\Delta$ bits of advice to encode the color of this cluster. If the cluster has a radius strictly smaller than $r/2-1$, then it means that the whole graph is contained in the cluster, and hence no advice is needed to compute a coloring. Hence, assume that the radius is at least $r/2-1$. Consider a BFS tree starting from the center $v$ of the cluster. By assumption, there exists a path that lies on the BFS tree, that does not contain $v$ and any of its neighbors, that does not contain nodes that have edges connected to nodes outside the cluster and all nodes at distance at most $\alpha$ from them, and that contains at least $r/2-4-\alpha$ nodes. By picking an $(\alpha,\alpha)$-ruling set on such a path, we obtain a set of size at least $(r/2-4-\alpha)/(2\alpha+1)=(50\alpha^{2}\log\Delta-4-\alpha)/(2\alpha+1)>10\alpha\log\Delta$ vertices that can be used in our advice schema. We sort these nodes by their IDs, and we assign to them, in order, the bits of the color of the cluster.
Clusters with degree larger than $\Delta^{10\alpha}$
By Lemma 5 we know that all clusters of degree larger than $\Delta^{10\alpha}$ have at least $\Delta^{9\alpha-2}$ internal vertices that are $\alpha$ hops away from their border. They either had it in the clustering around $I$, or they are clusters that survived the refining process and Lemma 5 guarantees those internal vertices exist.
In such a large set of vertices, there exists an $\alpha$-independent set of size at least $\Delta^{9\alpha-2}/\Delta^{\alpha+1}>\Delta^{7\alpha}$, for $\Delta>1$. Such an independent set can be found e.g. by a greedy algorithm. Adding a constraint that vertices cannot be too close to the center of the cluster takes away $1$ vertex from the set (the greedy algorithm used for construction can start from choosing the center). Therefore, we have at least $\Delta^{7\alpha}$ vertices, which for $\Delta>1$ and sufficiently large $\alpha$ is larger than $250\alpha^{2}\log^{2}\Delta+$ 2.
Decoding algorithm
Firstly, the algorithm identifies the clustering – all vertices with advice indicating that they belong to $I$ initiate a BFS computation, each vertex $v$ joins cluster of the initiator $i$ of the first BFS that reached $v$ (breaking ties in favor of the smallest neighbor ID). This gives the initial partition into clusters. Then, the clusters act upon the advice denoting whether they are in $I^{\prime}$ if they do not belong to $I^{\prime}$, their vertices join the closest high-degree cluster.
@@
Finally, as mentioned, to compute a coloring of vertices, each cluster-center can compute a coloring of the vertices in its cluster (it knows the topology of the cluster), and broadcast the computed colors to all vertices in the cluster. Overall, the time needed for computation is proportional to the radius of the clusters, which is $O(\alpha^{2}\log\Delta)$. ∎
6.2. $\Delta$-coloring
@@
The remaining part of Section 6.2 is dedicated to the proof of Lemma 6.
The algorithm in our schema relies on ideas similar to those used in (Panconesi and Srinivasan, 1992). On a somewhat abstract level, both algorithms start with uncoloring all vertices with color $\Delta+1$. Then, both have two steps - first they perform some local computations that reduce the set uncolored vertices to a set of uncolored vertices that are far apart. Then, given that the uncolored vertices are far apart, they can be simultaneously handled in a fairly simple way. The difference between our algorithm and that of (Panconesi and Srinivasan, 1992) is in what is considered far apart and what can be handled in a fairly simple way.
In the remaining part of the proof of Lemma 6, we first recall parts of the algorithm from (Panconesi and Srinivasan, 1992), and only afterwards, we describe our advice schema. Even though we do not implement the algorithm from (Panconesi and Srinivasan, 1992) directly, we recall it first as we use some building blocks of that algorithm to define the behavior of our algorithm and claim that it is indeed correct.
Some details of distributed $\Delta$-coloring algorithm (Panconesi and Srinivasan, 1992)
In Lemma 7, we formulate the observation from (Panconesi and Srinivasan, 1992) that is the most relevant for designing our $\Delta$-coloring advice schema.
Lemma 0.
@@
  • vertices from $X$ either have a degree smaller than $\Delta$ or have two neighbors with the same color in the given partial coloring, none on the path shortest path from $I$ to $X$,
@@
The set $I$ from Lemma 7 in the informal description above is the set of vertices that are far apart. Lemma 7 essentially says that if we can reduce the problem of extending partial $\Delta$-coloring (with no uncolored neighboring vertices) to a variant of the same problem where any two uncolored vertices are at least $2c\log_{\Delta}n$ hops from each other, then it can be solved in additional $c\log_{\Delta}n$ rounds. Then, the authors of (Panconesi and Srinivasan, 1992) provide the required reduction, which concludes their $\Delta$-coloring algorithm.
@@
The reduction starts with choosing as $\Omega(\log_{\Delta}n),O(\log^{2}_{\Delta}n)$ ruling set $R$, on the set of uncolored vertices (both vertices and distance constraints are held only on uncolored vertices, but the distance is defined by shortest paths using all edges). Then, each vertex finds a shortest path to the closest vertex from set $R^{\prime}$, which is $R$ extended by vertices of degree smaller than $\Delta$. Now, using these shortest path trees, the algorithm recolors vertices, considering vertices in a leaf-to-root order. Whenever an uncolored vertex is considered, it uncolors its parent, and then chooses a color from the palette of colors not used by its neighbors.
@@
The technical problem here is that neighboring uncolored vertices may need to choose color simultaneously, and as such some symmetry breaking is needed. To that end, the authors of (Panconesi and Srinivasan, 1992) use a randomized approach, that in expected $O(\log n)$ rounds can process one layer of the trees. The authors of (Ghaffari et al., 2021) handle an analogous problem by a slightly different approach, and employ Theorem 8 to handle that issue.
Theorem 8.
[…]
Our advice schema
@@
While it’s possible to design a schema for the 2 in a direct way, considering the problems of recoloring vertices from $I$ and all other uncolored vertices separately allows us to focus on one difficulty at a time. For each of the two subproblems we design a composable advice schema, and the final schema from Lemma 6 is obtained by composing them using Lemma 1.
@@
First, we show the part of our algorithm that reduces the general problem to the problem of coloring vertices from some $\Omega(\log_{\Delta}n)$ independent set $I$. To recall, in order to handle this subproblem, the algorithm from (Panconesi and Srinivasan, 1992) solves a sequence of $O(\log^{2}n)$ list coloring problems. By a sequence of $O(\log^{2}n)$ problems, we mean that we can’t solve them simultaneously, as input for some of those problems depends on the output of other problems, and those dependencies can form chains of length $O(\log^{2}n)$. We leverage advice in a way, that allows us to solve some of those problems simultaneously, by encoding the answers to a subset of those problems, in a way that allows running more of the list coloring algorithms in parallel.
Problem 3.
@@
We are given a $\Delta$ colorable graph $G$ with maximum degree $\Delta$, with a proper $\Delta+1$ coloring $C$. The goal is to compute a proper partial $\Delta$-coloring of $G$, in which any uncolored vertices are at a distance at least $2c\log_{\Delta}n$, where $c$ is the constant $c$ from Lemma 7.
Lemma 0.
@@
Let $\mathcal{G}_{\Delta}$ be the set of all $\Delta$-colorable graphs with vertices colored according to some proper $\Delta+1$ coloring. Let $\Pi_{\Delta}$ be the partial $\Delta$-vertex coloring problem described in 3. Then, there exists a $(\mathcal{G}_{\Delta},\Pi_{\Delta},\gamma_{0},A,T)$-composable advice schema, where $\gamma_{0}=O(1)$, $A(c,\gamma)=\Theta(\gamma^{3})$, and $T(\alpha,\Delta)=O(\alpha\log\Delta\cdot\sqrt{\Delta\log\Delta})$.
Proof.
@@
The general idea is to consider the same shortest path spanning forest used in (Panconesi and Srinivasan, 1992). However, instead of processing the trees directly, we choose a set of relay vertices, and regular uncolored vertices use them to define another set of shortest path trees, later used for recoloring.
A proper set of relay vertices can be found as follows. Let us consider an $\Omega(\log_{\Delta}n),O(\log^{2}_{\Delta}n)$ ruling set $R$, and the shortest path spanning forest $F$ as in the distributed algorithm (Panconesi and Srinivasan, 1992). To define the relay vertices, we consider vertices in $F$ in layers.
@@
  • for each relay vertex there is some space to deposit advice bits
  • relay vertices from layer $i$ are not too far from the relay vertices from layer $i-1$
  • each uncolored vertex is close to a relay vertex

### v2

@@ Appendix C $\Delta$-coloring of $\Delta$-colorable graphs
The remaining part of this section is dedicated to proving Theorem˜C.1. The formal definition of advice schema and composability are presented in Definition˜3.2 and Definition˜3.4, respectively. To recall, in order to prove that a composable schema exists, for any given $\gamma\geq\gamma_{0}$ (and hence in our case a large-enough constant), and for any given $\alpha\geq A(c,\gamma)$, we need to provide a schema of distributing advice bits to the vertices of the graph such that, in any ball of radius $\alpha$ there are at most $\gamma_{0}$ bit-holding nodes, and bit-holding nodes must have at most $c\alpha/\gamma^{3}$ bits each. Moreover, given such a bit assignment, there must exist an algorithm that can solve $\Delta$-coloring in $\max(O(\alpha^{2}\log\Delta),O(\alpha\log\Delta\cdot\sqrt{\Delta\log\Delta})$ time.
@@
  2. 2.
    Then, the algorithm reduces the number of colors to $\Delta+1$, by using a standard distributed coloring algorithm, which runs in $O(\sqrt{\Delta\log\Delta})$ rounds when an initial $O(\Delta^{2})$-coloring is given [24, 6, 46].
  3. 3.
    Finally, the algorithm reduces the number of colors from $\Delta+1$ to $\Delta$, inspired by distributed algorithms from [53, 31], leveraging the advice bits to terminate in just $O(\log\Delta\cdot\sqrt{\Delta\log\Delta})$ rounds.
The second step is a simple application of distributed algorithms with no advice, while the first and third algorithms rely on some given advice. Below, we formally state the problems to be solved in steps 1 and 3, and then we design composable schemas for them. Later, we will compose the schemas into a single one by applying Lemma˜E.1, obtaining Theorem˜C.1.
C.1 $O(\Delta^{2})$-vertex coloring
@@
Problem C.3 ($O(\Delta^{2})$-coloring).
@@
The algorithm roughly follows the approach that can be used to color a graph via a network decomposition, which can be described as follows. Let us consider a partition of vertices into disjoint clusters, each cluster having only $\Delta^{O(\alpha^{2}\log\Delta)}$ adjacent clusters and a small diameter. In order to find a proper $\Delta^{O(\alpha^{2}\log\Delta)}$ coloring, we can use a proper $\Delta+1$ coloring of vertices in each cluster, and a proper $\Delta^{O(\alpha^{2}\log\Delta)}$ coloring of the cluster graph (i.e., a coloring of the graph where each cluster is contracted to a point, or, in other words, an assignment of colors to clusters such that adjacent clusters have different colors). The color of a vertex $v$ in a cluster $C$ is the tuple ((color of $v$ in $C$, color of $C)$. The first part of the tuple ensures that neighboring vertices in the same cluster have different colors, and the second part ensures that vertices in two adjacent clusters have different colors. The total number of colors we use is still $\Delta^{O(\alpha^{2}\log\Delta)}$. The number of colors we use can be then reduced to $O(\Delta^{2})$ by using a few rounds of Linial’s coloring algorithm [45].
Lemma C.6.
@@
[45] Let us consider a graph with maximum degree $\Delta$ and some coloring of its vertices with colors $1,2,\dots,c$. There exists $2$ round distributed LOCAL algorithm that finds a proper $O(\Delta^{2})$ coloring, if $c\in\operatorname{poly}(\Delta)$, or an $O(\Delta^{2}\log c)$ coloring otherwise.
@@
Given that the total number of colors before we start using the reduction described in Lemma˜C.6 is $\Delta^{O(\alpha^{2}\log\Delta)}$, we need $O(\log^{*}\alpha)$ rounds to reduce the number of colors down to $O(\Delta^{2})$.
@@
In the remaining part of the proof of Lemma˜C.4 we show that we can define an adequate clustering, a distribution of advice bits, and an algorithm that using these advice bits finds a proper $\Delta^{O(\alpha^{2}\log\Delta)}$ coloring of the cluster graph.
@@
In the following, the degree of a cluster is the number of edges with exactly one endpoint in the cluster. The idea of the advice schema of this section is similar to the idea behind the schema for graphs with sub-exponential growth. Essentially, we show that we can define a clustering with clusters that have sufficiently many internal vertices to encode the color of the cluster. However, to provide such a clustering, we rely on a slightly different idea, namely: we split all clusters into $O(\alpha^{2}\log\Delta)$ buckets, and let the $i$th bucket contain every cluster whose degree is in the interval $\Delta^{i-1},\Delta^{i})$. Each bucket receives its own disjoint palette of colors; the $i$th bucket receives a palette of size $\Delta^{i}$. The clusters from the $i$th bucket will be colored only with those colors, and as such the number of bits describing such a color is at most $(i+1)\log\Delta$. For any $\Delta>1$, the total number of colors in bucket $i$ and all smaller buckets is $\sum_{j=1}^{i}\Delta^{j}<\Delta^{i+1}$.
@@
To define the clustering, we proceed as follows. Let $r=100\alpha^{2}\log\Delta$. We pick an $(r,r)$-ruling set $I$ from $G$. Then, we assign each vertex from $G$ to the closest vertex from $I$ (breaking ties in an arbitrary consistent manner, e.g., by joining the cluster of the smallest ID neighbor). That concludes the first part of our construction. Currently, we have clusters satisfying that all nodes that are within distance $r/2-1=50\alpha^{2}\log\Delta-1$ from a cluster center $v$ are guaranteed to belong to the cluster of $v$, and such that the degree of the clusters is bounded from above by $\Delta^{r+1}$. However, for our advice schema, we need the clusters with degrees larger than $\Delta^{10\alpha}$ to have at least $100\alpha^{2}\log^{2}\Delta+2$ internal vertices that are at least $\alpha$ hops away from the border of the cluster.
@@
We deem a cluster to be broken when it has degree at least $\Delta^{10\alpha}$ but does not have at least $\Delta^{9\alpha-2}$ vertices of distance at least $\alpha$ from its border. In the refining process, we consider all broken clusters one by one, and we ‘repair’ each broken cluster by destroying its neighbors, and possibly destroying other nearby clusters. Whenever we consider a broken cluster $B$ that is still in $C$, we remove from $C$ all other clusters ($\neq B$) that contain at least one vertex whose distance from the border of $B$ is smaller than $2\alpha+1$. Thus, the set $I^{\prime}$ still contains the center $i_{B}$ of said broken cluster, but has lost one or more other cluster centers.
@@
Proof C.8.
For any broken cluster $B$ that is not removed from $C$, the refining process removed all clusters (in particular all high-degree clusters) with vertices at a distance at most $2\alpha+2$ from the border of $B$. As such, the re-partition step assigns to $B$ all vertices that are at distance at most $\alpha$ from B’s border, as any other cluster is at least $\alpha+1$ hops away, and $B$ is at most $\alpha$ hops away. Thus, all vertices that were originally in $B$ are now at least $\alpha$ hops away from their new border. In conclusion, $B$ contains at least $\Delta^{10\alpha}$ nodes that are at least $\alpha$ hops away from their new border.
@@
Clusters with degree at most $\Delta^{10\alpha}$.
If the degree of a cluster is smaller than $\Delta^{10\alpha}$, we need $10\alpha\log\Delta$ bits of advice to encode the color of this cluster. If the cluster has a radius strictly smaller than r/21r/2-1, then it means that the whole graph is contained in the cluster, and hence no advice is needed to compute a coloring. Hence, assume that the radius is at least r/21r/2-1. Consider a BFS tree starting from the center $v$ of the cluster. By assumption, there exists a path that lies on the BFS tree, that does not contain $v$ and any of its neighbors, that does not contain nodes that have edges connected to nodes outside the cluster and all nodes at distance at most $\alpha$ from them, and that contains at least $r/2-4-\alpha$ nodes. By picking an $(\alpha,\alpha)$-ruling set on such a path, we obtain a set of size at least $(r/2-4-\alpha)/(2\alpha+1)=(50\alpha^{2}\log\Delta-4-\alpha)/(2\alpha+1)>10\alpha\log\Delta$ vertices that can be used in our advice schema. We sort these nodes by their IDs, and we assign to them, in order, the bits of the color of the cluster.
Clusters with degree larger than $\Delta^{10\alpha}$.
By Lemma˜C.7 we know that all clusters of degree larger than $\Delta^{10\alpha}$ have at least $\Delta^{9\alpha-2}$ internal vertices that are $\alpha$ hops away from their border. They either had it in the clustering around $I$, or they are clusters that survived the refining process and Lemma˜C.7 guarantees those internal vertices exist.
In such a large set of vertices, there exists an $\alpha$-independent set of size at least $\Delta^{9\alpha-2}/\Delta^{\alpha+1}\allowbreak>\Delta^{7\alpha}$, for $\Delta>1$. Such an independent set can be found e.g. by a greedy algorithm. Adding a constraint that vertices cannot be too close to the center of the cluster takes away $1$ vertex from the set (the greedy algorithm used for construction can start from choosing the center). Therefore, we have at least $\Delta^{7\alpha}$ vertices, which for $\Delta>1$ and sufficiently large $\alpha$ is larger than $50\alpha^{2}\log^{2}\Delta+2$.
Decoding algorithm.
Firstly, the algorithm identifies the clustering – all vertices with advice indicating that they belong to $I$ initiate a BFS computation, each vertex $v$ joins cluster of the initiator $i$ of the first BFS that reached $v$ (breaking ties in favor of the smallest neighbor ID). This gives the initial partition into clusters. Then, the clusters act upon the advice denoting whether they are in $I^{\prime}$ - if they do not belong to $I^{\prime}$, their vertices join the closest high-degree cluster.
@@
Finally, as mentioned, to compute a coloring of vertices, each cluster-center can compute a coloring of the vertices in its cluster (it knows the topology of the cluster), and broadcast the computed colors to all vertices in the cluster. Overall, the time needed for computation is proportional to the radius of the clusters, which is $O(\alpha^{2}\log\Delta)$.
C.2 $\Delta$-coloring
@@
The remaining part of Section˜C.2 is dedicated to the proof of Lemma˜C.10.
The algorithm in our schema relies on ideas similar to those used in [53]. On a somewhat abstract level, both algorithms start with uncoloring all vertices with color $\Delta+1$. Then, both have two steps - first they perform some local computations that reduce the set uncolored vertices to a set of uncolored vertices that are far apart. Then, given that the uncolored vertices are far apart, they can be simultaneously handled in a fairly simple way. The difference between our algorithm and that of [53] is in what is considered far apart and what can be handled in a fairly simple way.
In the remaining part of the proof of Lemma˜C.10, we first recall parts of the algorithm from [53], and only afterwards, we describe our advice schema. Even though we do not implement the algorithm from [53] directly, we recall it first as we use some building blocks of that algorithm to define the behavior of our algorithm and claim that it is indeed correct.
Some details of distributed $\Delta$-coloring algorithm [53].
In Lemma˜C.11, we formulate the observation from [53] that is the most relevant for designing our $\Delta$-coloring advice schema.
Lemma C.11.
@@
  • vertices from $X$ either have a degree smaller than $\Delta$ or have two neighbors with the same color in the given partial coloring, none on the shortest path from $I$ to $X$,
@@
The set $I$ from Lemma˜C.11 in the informal description above is the set of vertices that are far apart. Lemma˜C.11 essentially says that if we can reduce the problem of extending partial $\Delta$-coloring (with no uncolored neighboring vertices) to a variant of the same problem where any two uncolored vertices are at least $2c\log_{\Delta}n$ hops from each other, then it can be solved in additional $c\log_{\Delta}n$ rounds. Then, the authors of [53] provide the required reduction, which concludes their $\Delta$-coloring algorithm.
@@
The reduction starts with choosing an $\Omega(\log_{\Delta}n),O(\log^{2}_{\Delta}n)$ ruling set $R$, on the set of uncolored vertices (both vertices and distance constraints are held only on uncolored vertices, but the distance is defined by shortest paths using all edges). Then, each vertex finds a shortest path to the closest vertex from set $R^{\prime}$, which is $R$ extended by vertices of degree smaller than $\Delta$. Now, using these shortest path trees, the algorithm recolors vertices, considering vertices in a leaf-to-root order. Whenever an uncolored vertex is considered, it uncolors its parent, and then chooses a color from the palette of colors not used by its neighbors.
@@
The technical problem here is that neighboring uncolored vertices may need to choose color simultaneously, and as such some symmetry breaking is needed. To that end, the authors of [53] use a randomized approach, that in expected $O(\log n)$ rounds can process one layer of the trees. The authors of [31] handle an analogous problem by a slightly different approach, and employ Theorem˜C.12 to handle that issue.
Theorem C.12 ([24, 6, 46]).
There is a deterministic distributed algorithm that given proper $O(\Delta^{2})$ vertex coloring solves the $({\deg}+1)$-list coloring problem in time $O(\sqrt{\Delta\log\Delta})$.
Our advice schema.
@@
While it’s possible to design a schema for the ˜C.9 in a direct way, considering the problems of recoloring vertices from $I$ and all other uncolored vertices separately allows us to focus on one difficulty at a time. For each of the two subproblems we design a composable advice schema, and the final schema from Lemma˜C.10 is obtained by composing them using Lemma˜E.1.
@@
First, we show the part of our algorithm that reduces the general problem to the problem of coloring vertices from some $\Omega(\log_{\Delta}n)$ independent set $I$. To recall, in order to handle this subproblem, the algorithm from [53] solves a sequence of $O(\log^{2}n)$ list coloring problems. By a sequence of $O(\log^{2}n)$ problems, we mean that we cannot solve them simultaneously, as input for some of those problems depends on the output of other problems, and those dependencies can form chains of length $O(\log^{2}n)$. We leverage advice in a way, that allows us to solve some of those problems simultaneously, by encoding the answers to a subset of those problems, in a way that allows running more of the list coloring algorithms in parallel.
Problem C.13.
@@
We are given a $\Delta$ colorable graph $G$ with maximum degree $\Delta$, with a proper $\Delta+1$ coloring $C$. The goal is to compute a proper partial $\Delta$-coloring of $G$, in which any uncolored vertices are at a distance at least $2c\log_{\Delta}n$, where $c$ is the constant $c$ from Lemma˜C.11.
Lemma C.14.
@@
Let $\mathcal{G}_{\Delta}$ be the set of all $\Delta$-colorable graphs with vertices colored according to some proper $\Delta+1$ coloring. Let $\Pi_{\Delta}$ be the partial $\Delta$-vertex coloring problem described in ˜C.13. Then, there exists a $(\mathcal{G}_{\Delta},\Pi_{\Delta},\gamma_{0},A,T)$-composable advice schema, where $\gamma_{0}=O(1)$, $A(c,\gamma)=\Theta(\gamma^{3})$, and $T(\alpha,\Delta)=O(\alpha\log\Delta\cdot\sqrt{\Delta\log\Delta})$.
Proof C.15.
@@
The general idea is to consider the same shortest path spanning forest used in [53]. However, instead of processing the trees directly, we choose a set of relay vertices, and regular uncolored vertices use them to define another set of shortest path trees, later used for recoloring.
A proper set of relay vertices can be found as follows. Let us consider an $\Omega(\log_{\Delta}n)$, $O(\log^{2}_{\Delta}n)$ ruling set $R$, and the shortest path spanning forest $F$ as in the distributed algorithm [53]. To define the relay vertices, we consider vertices in $F$ in layers.
@@
  • for each relay vertex there is some space to deposit advice bits,
  • relay vertices from layer $i$ are not too far from the relay vertices from layer i1i-1,
  • each uncolored vertex is close to a relay vertex.